Find the range of values of a for which the following function is periodic with period of 4.
sin[a]x
[] is an integral part function.

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
f(x) = sin[a]x
Period of function:
\frac{2\pi }{[a]}
Given:
\frac{2\pi }{[a]} = 4
[a] = \frac{\pi }{2}
There is no value of a for which above equation satisfy.
a\in \phi